Calculate the occupancy rate for Gila Monster Memorial Medical Center, given the following data:

Inpatient days = 55569

Outpatient visits = 131728

Number of licensed beds = 29

Operating rooms = 4

occupancy rate=
Answer #1

Occupancy Rate = Inpatient Days of Care / Bed Days Available x 100

Occupancy Rate = 55,569 / (29 x 365) x 100

Occupancy Rate = 55,569 / 10,585 x 100 = 524.98% occupancy.

Note: There seems to be something off about the information in the question. Either the number of beds count is wrong or the number of years (assumed 1 year in solution) for which the Inpatient data is given is incorrect.

Generally Outpatient visits and operating rooms are not considered for computing the occupancy rate of a hospital.